A circle has the equation x^2+y^2+3x-4y+4=0 graph circle using center (h,k)and radius r
Answer by ewatrrr(24785)

 
Hi,
x^2+y^2+3x-4y+4=0
x^2+3x + 9/4 + y^2-4y +4= -4 + 9/4 + 4
(x+3/2)^2 + (y-2)^2 = 9/4 , C(-3/2,2) r = 3/2